Will AgIO3 precipitate when 20 mL of 0.010M AgIO3 is mixed with 10 mL of 0.015M NaIO3? Ksp of AgIO3 is 3.1× 10-8

1
Expert's answer
2022-04-25T13:36:04-0400

Will AgIO3 precipitate when 20 mL of 0.010 M AgNO3 is mixed with 10 mL of 0.015 M NaIO3?

Ksp of AgIO3 is 3.1×10−8


Solution:

AgNO3 → Ag+ + NO3

So, the molarity of Ag+ is same as the molarity of AgNO3 (0.010 M)

NaIO3 → Na+ + IO3

So, the molarity of IO3 is same as the molarity of NaIO3 (0.015 M)


Moles of Ag+ = (0.010 M) × (20 mL) × (1L / 1000 mL) = 0.0002 mol

Moles of IO3 = (0.015 M) × (10 mL) × (1L / 1000 mL) = 0.00015 mol


Total volume = 20 mL + 10 mL = 30 mL = 0.03 L

Therefore,

[Ag+] = (0.0002 mol) / (0.03 L) = 0.00667 M

[IO3] = (0.00015 mol) / (0.03 L) = 0.0050 M


AgIO3 → Ag+ + IO3

Qsp = [Ag+] × [IO3] = (0.00667) × (0.0050) = 3.33×10−5


Since Qsp > Ksp

3.33×10−5 > 3.1×10−8

Therefore,

AgIO3 will precipitate


Answer: AgIO3 will precipitate